Calculate the income elasticity of commodity A if income of the consumer changes from $1500 to $3400 and the quantity also changes from 150 to 405

1
Expert's answer
2022-02-17T04:38:24-0500

The income elasticity of commodity A is:

"Ei = \frac{405 - 150} {3400-1500} ×\frac{3400+1500} {405+150} = 1.18."

So, the commodity A is a normal good.